OpenStudy (bibby):

Let X be a RV with cumulative density function \(f(x) = x^2\) for \(0 \le x \le 1\)

OpenStudy (bibby):

Show that this is a valid cumulative density function. (Two conditions must be checked). would that mean show that \(\int\limits_{- \infty}^{\infty}f(x)=1\) and that \(f(x)\ge 0\)?

OpenStudy (bibby):

while I'm here, part b is What is the probability that X is between 1 and 3?

OpenStudy (freckles):

A) For all a and b, if a<b then f(a) <=f(b) (f is non-decreasing) B) x->-inf implies f(x)->0 C) x->inf implies f(x)->1 D) f is right continuous --- you know i think that 1 thing comes from if the cdf is a pdf

OpenStudy (freckles):

i'm about your integral being 1

OpenStudy (freckles):

I think \[P(1\le X \le 3)=f(3)-f(1)\]

OpenStudy (freckles):

but hey f(x)=x^2 is only defined for 0<=x<=1 ?

OpenStudy (freckles):

what is f when x is 3?

OpenStudy (bibby):

I think it's p(X=x)

OpenStudy (bibby):

uhh 9?

OpenStudy (freckles):

so f(x)=x^2 when 0<=x<=?

OpenStudy (freckles):

I'm just confused because it looks like you only have f(x)=x^2 for when 0<=x<1

OpenStudy (bibby):

f(1)-f(0) = 1-0=1

OpenStudy (freckles):

<=*

OpenStudy (bibby):

I'm confused as well, I assume my professor is going to clear it up later tonight (hopefully)

OpenStudy (freckles):

I can show you an example that makes more sense to me And I'm saying it makes more since because we have a function defined for where we do want to find the probability for certain values for x. let f(x)=0 if x<0 =x/2 if 0<=x<1 =1 if 1<=x --------------------- \[P(-1<X \le \frac{1}{2})=f(\frac{1}{2})-f(-1)=\frac{\frac{1}{2}}{2}-0=\frac{1}{4}-0=\frac{1}{4}\]

OpenStudy (freckles):

maybe your f(x) was suppose to be f(x)=0 if x<0 =x^2 if 0<=x<1 =1 of 1<=x

OpenStudy (freckles):

If that is so then \[P(1 \le X \le 3)=f(3)-f(1)=1-1=0\]

OpenStudy (freckles):

which sorta makes sense there is a 0 probability of it happening since your function isn't even defined there

OpenStudy (bibby):

That explains a lot, thank you. Off to memorize some distributions now, thanks for the help :)

OpenStudy (freckles):

oh we could have wrote f as f(x)=0 if x<0 =x^2 if 0<=x<=1 =1 if x>1 you know since you had that in your initial post (that middle inequality) but it doesn't matter since x^2=1 when x=1

OpenStudy (freckles):

np good luck

Parth (parthkohli):

@bibby dude.